# ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries

### v2

@@ 1 Introduction
[ht!] public RowBuilder int16(String name){ ColumnInt16 column = new ColumnInt16(_columns.size(), name, _offset); _offset += column.length(); _columns.add(column); return this;} Summary: ……. This method is used to add a new column of data type int16 (16-bit integer) to the existing data structure. It creates a new ColumnInt16 object with the given name and size (16 bits), updates the offset value to accommodate the new column, and adds the new column to the _columns list……… Create a new ColumnInt16 object with the given name and size (16 bits). ……………. Example 1: Confused data type in LLama3-70B Summary
@@
In this work, we study different factors that can lead to hallucination and list down a taxonomy to map the common causes easily. We also notice a lack of datasets to reliably research this topic. Therefore, we create a first-of-its-kind dataset for studying hallucination in code summarization with 411 summaries generated by seven different large language models, broken into 9933 entity-level samples. This dataset consists of code and a corresponding summary describing the code. The annotation consists of: a) Named Entity Recognition b) Entity Level Description Verification (c) Overall Summary Quality based on inaccuracies (not focusing on code completeness or conciseness). We then introduce a framework that focuses on validating the correctness of the generated summary with respect to code entities. For this, we verify if the entities discussed in the summary are present in the code and correctly described in the summary. The framework leverages code parsers like javalang [^1] https://github.com/c2nes/javalang to list the different entities in the code snippet and prompt-based approaches to detect entities in summary. We note that detecting entities in the generated summary is more difficult due to the high degree of polysemy Tabassum et al. (2020). For example, entities like "list", "while", "if", etc, can be a code entity or natural language entities. This necessitates our reliance on large language models with high reasoning capabilities for detecting entities on the summary side. We then map the detected entities from the summary to code by using string-matching heuristics. The sentences with unmapped entities can be considered as ungrounded (source of extrinsic hallucination). For each mapped entity, we then have a tuple <code, entity, intent-related sentence>, where the intent-related sentence can be considered as the sentence in summary mentioning the entity. The final step is to verify each tuple from the summary for intrinsic hallucination and aggregate them to calculate the overall correctness of the code summary. Our experiments demonstrate the importance of localizing entities in the summary for effective hallucination detection. Our contributions are :
@@ 2 Related Work
Hallucination in Natural Langauge: Rawte et al. (2024); Sahoo et al. (2024) review recent advances in hallucination detection in natural language, emphasizing its practical significance. Recently, prompt-based methods (Arora et al. (2022), Manakul et al. (2023), Agrawal et al. (2023), Dhuliawala et al. (2023)) are being used to detect hallucinations in the text produced by LLMs. Xiao and Carenini (2022) and Zhang et al. (2022) attempt to address entity-level verification in natural language inputs. Both of these works involve improving the correctness of natural language summaries and do not discuss anything in the context of code. Maynez et al., 2020; Ji et al., 2023 discuss further fine-graining of hallucination in natural language as intrinsic and extrinsic hallucination. More specifically, Intrinsic hallucination occurs when the given text contradicts the reference, while Extrinsic hallucination happens when the text cannot be verified against the reference. We use a similar convention in our paper.
Hallucination in Code Generation: The code generation space has captured significant attention due to its practical significance in software development. Jiang et al. (2024) discusses recent developments in code generation and suggests the importance of addressing hallucination for improving the reliability of LLMs. Liu et al. (2024) studies hallucination in code generation and proposes a categorization that encompasses five categories of hallucinations based on the conflicting objectives and varying degrees of deviation observed in code generation. Tian et al., 2024; Agarwal et al., 2024; Spracklen et al., 2024 further advanced the field by proposing valuable datasets and frameworks to tackle hallucination in code generation. These studies indicate the tendencies of LLMs to produce code that is syntactically correct and even semantically reasonable but fails to execute as intended or meet the specified requirements.
Although there have been significant advances in hallucination detection in code generation, the field of code summarization is still nascent. Recently Kang et al., 2024; Zhang, 2024 have investigated inconsistencies in comment generation, but their focus has been limited to specific aspects such as verifying design constraints like parameter types and ranges. Additionally, some of their approaches, such as generating test cases for comments, face practical challenges due to heavy reliance on execution environments. In contrast, our approach aims to validate the entire functionality described in the output, independent of external environments or dependencies. By grounding entities based on the input code and verifying their intent, our framework offers a more reliable and less dependent solution.
@@ 3 Datasets
To create the hallucination dataset for code summarization, we consider code snippets from Java programming language and CodeXGLUE Lu et al. (2021) – Code-To-Text dataset. We focused on Java programming language due to its widespread relevance in the industry. It offers a rich set of entities (such as classes, methods, and variables) due to its structured design and strict typing system. We report the statistics for the data in Table 1 and describe the curation below:
Summary Generation: We generate summaries from CodeXGLUE by prompting seven different LLMs (Appendix A) on 600 different code snippets and consider $\sim$10of this data for the final hallucination annotation task (Table 1). By generating multiple summary variants, we can assess the extent of hallucination generation by different LLMs and study the hallucination detection techniques under diverse conditions.
@@
Hallucination Labeling: For each detected code entity in a summary, all sentences describing that entity are considered relevant. To account for the scenario where the relevant sentence can be noisy, we introduced a third label, "IRRELEVANT", which can be used to evaluate the performance of the intent-detection module and removed during the preprocessing. Thus, we obtain tuples of (code, entity, relevant sentences) for each entity. A total of 9933 such tuples, sampled from 441 summaries, were selected for human annotation (hired based on volunteering) to detect hallucinations. Out of these, 4354 tuples (from 222 summaries) were independently reviewed by two different sets of annotators, leading to a Cohen Kappa score of 0.72, implying high agreement. The conflicts were resolved by two independent meta-annotators. Annotators were tasked to (i) correct named entities by removing incorrect ones, adding missing ones, and adjusting entity types; (ii) evaluate if each sentence accurately describes the entity in the code snippet, marking the entity-sentence pair as correct, incorrect, or irrelevant; and (iii) rate the summary based on hallucination severity as "good," "fair," or "poor." We utilize these labels to establish the definition of summary-level hallucination assessment derived from entity-level hallucination. We observe that, on average, 1.33 entities were marked as hallucinated for the summaries rated as ’fair" or ’poor". Therefore, we consider a summary as hallucinated if at least one of the entities is hallucinated. After pre-processing, we consider the instance with labels "CORRECT" and "INCORRECT" in human data and treat the "IRRELEVANT" label predicted by the model as "INCORRECT".
@@ 4 Categorization of Factors for Hallucination in Code Summarization
In this section, we describe the various factors that could lead to hallucination in code summaries (Figure 2) based on what we learned from the annotation process. This classification, based on the underlying factors of hallucination, offers insights into the generative behaviors of language and code models. We categorize hallucination factors in code summarization into the following groups:
@@
1) Contextual code This occurs when the model is unable to correctly explain the code present in the input. This may happen if the model hasn’t encountered the code’s functionality during training or if it’s dealing with a low-resource language like COBOL, where basic rules might be misrepresented in the summary.
@@
HC3: Code Complexity: This category highlights the model’s tendency to produce incorrect code summaries due to high code complexity. From another perspective, the model may not have enough reasoning capabilities to understand the code correctly and may deviate from the instruction given by the user. These complexities can arise due to various reasons, such as 1) Length: Longer code tends to be more complex because it often requires more understanding, has more potential points of failure, and may involve more interdependencies. 2) Lexical Complexity: This refers to scenarios where complex vocabulary in code, including diverse operands (variables, constants) and operations (functions, operators), increases complexity due to the higher number of elements to track and understand. 3) Logical complexity: This refers to scenarios where the code’s logic is complex, often due to high cyclomatic complexity, indicating many independent paths. Complexity increases with numerous paths or method invocations, especially when these are not directly visible or are called from distant parts of the codebase.
@@
HC4: Natural language Context: This category refers to cases where natural language in code snippets, such as outdated comments or log statements, causes hallucinations in code summaries. In the code snippet shown in HC4 Example, the LLama3-70B model incorrectly infers that the property variable contains a list of key-value pairs inferred from a commented line. However, the ‘property" variable contains an alphanumeric string followed by one or more semi-colons.
@@ 5.1 Entity Verification
Entity Extraction from summary: Code summaries often focus on properties like Variables, Classes, Functions, and Tables, which cannot be categorized using traditional NLP tagsets. To address this, Tabassum et al. (2020) propose the task of entity detection in code summaries and introduce a relevant NER tagset that includes categories such as CLASS, VARIABLE, FUNCTION, LIBRARY, etc. We adopt this tagset for extracting entities from code summaries (Prompt: Appendix A Figure 3). Since we are leveraging LLMs to recognize the entities, there is a possibility of encountering hallucinations during this step. This involves the scenario when the model fabricates some entities from summaries and outputs the entities that might not be present in the code summary itself. To address this, we add a filtration step which removes any entities not present in the summary. We evaluate Gemini and GPT-4-Omni for the task of Code NER on the collected human data and report their results in Appendix C. We also evaluate the open-source model for this task as an additional contribution. We observe a strong correlation between GPT-4-Omni predictions and human data, making it suitable for entity detection in our framework.
@@ 5.2 Entity-Intent Verification
The presence of an entity in both the summary and code indicates that the entity is valid but does not warrant the correctness of the context in which it is discussed. For example, in Figure 1 jobId is a correct entity, but the context of retrieving jobID from the database is incorrect. To address this problem, we propose verifying whether the intent of each mapped entity is accurately described in the summary. We extract all sentences containing the entity of interest from the summary to form its intent context. To identify these relevant sentences that describe an entity’s intent, we explored two approaches: (1) prompting a language model to find relevant sentences and (2) using string-matching heuristics. Qualitative assessment showed that rule-based heuristics were more effective and efficient than prompt-based methods, which suffered from hallucinations. Therefore, we relied on string-matching-heuristics for our framework. After identifying the entity and intent context, we use LLMs with zero-shot prompting to verify their correctness against the code (Prompt: Appendix 5).
@@ 6 Experiments and Results
For our experiments, we consider summaries from instruction-tuned versions of the SOTA code and language models, from IBM-Granite family (20B-instruct; 34B-instruct) Mishra et al. (2024), Llama3 family (8B-instruct and 70B-instruct) Touvron et al. (2023), CodeLlama family (7B and 34B) Roziere et al. (2023) and Mistral family (7B-instruct) Jiang et al. (2023). For intent verification, the use of LLMs to verify their own generation may be challenging due to inherent confirmation bias Xie et al. (2023); Feng et al. (2024), which enables the model to agree with their own generations. To avoid confirmation bias, we chose GPT4-Omni Achiam et al. (2023) and Gemini-1.5-Flash Team et al. (2023) for our entity-intent verification task. Hyperparameters and other details can be found in Appendix (B). We discuss the results from two different aspects of evaluation below:
Model P R F1 A
@@
Gemini-1.5-Direct 0.41 0.49 0.25 0.32
Gemini-1.5-ETF* 0.68 0.64 0.51 0.52
GPT4-Omni-Direct 0.48 0.50 0.28 0.33
GPT4-Omni-ETF* 0.72 0.74 0.73 0.75
@@
Gemini-1.5 0.55 0.70 0.55 0.84
GPT4-Omni 0.59 0.69 0.61 0.92
Table 2: Hallucination Detection results. We consider two evaluation aspects: 1) Instance Level, which aims to label the entire summary, and 2) Entity level, which labels individual entities in summaries. Here, P, R, F1 and A refer to the precision, recall and F1 score and accuracy, respectively (macro avg). The ETF (*) (entity tracing framework) indicates the proposed framework
@@ 7.1 Quantitative Analysis
As shown in Table 3, Granite-20B produced shorter summaries, while Llama3-70B generated longer ones. Other models had similar average lengths, reflecting varying elaboration due to differences in training methodologies. For entity mapping, we observe that Llama3-70B has the most mapped entities, indicating the tendency of the model to stay grounded. While Granite-20B has the most unmapped entities, which indicates its tendency to produce content which may not be directly related to the code leading to extrinsic hallucination.
@@ 7.2 Predictive Analysis
Our framework can be utilized to capture the ungrounded entities within the summary. In the example given below, the model tries to guess the summary based on the mentioned keyword "HubException" and discusses a non-existent Hub API. For this example, the unmapped entities are $\{^{\prime}squid:s1166^{\prime},^{\prime}Hub^{\prime},^{\prime}squid:s1172^{\prime},^{\prime}Hubdatabase^{\prime}\}$. We can see that the ungrounded entities like "Hub" are captured here, leading to more fine-grained and interpretable detection of hallucination.
[ht!] private List<Transaction> retrieveTransactions(String rowStatusCd) throws HubException {................} Summary: This method retrieves transactions using the “Hub" API based on the input parameters provided by the “Hub database". It interacts with several key entities, including “squid:s1166" and “squid:s1172", to gather the data…… Analysis: Hallucination and Unmapped Entities

### v3

@@ 1 Introduction
[ht!] public RowBuilder int16(String name){ ColumnInt16 column = new ColumnInt16(_columns.size(), name, _offset); _offset += column.length(); _columns.add(column); return this;} Summary: ……. This method is used to add a new column of data type int16 (16-bit integer) to the existing data structure. It creates a new ColumnInt16 object with the given name and size (16 bits), updates the offset value to accommodate the new column………… Example 1 (LLama3-70B): Confused Data Type
@@
In this work, we study different factors that can lead to hallucination and list down a taxonomy to map the common causes easily. Noting a lack of datasets to reliably research this topic. Therefore, we create a first-of-its-kind dataset for studying hallucination in code summarization with 411 summaries generated by seven different large language models, broken into 9933 entity-level samples. This dataset consists of code and a corresponding summary describing the code. The annotation consists of: a) NER b) Entity Description Verification (c) Overall Summary Quality (not focusing on code completeness or conciseness). We then introduce a framework that evaluates the correctness of the generated summary. For this, we verify if the entities discussed in the summary are present in the code and correctly described in the summary. The framework leverages code parsers like javalang [^1] https://github.com/c2nes/javalang to list the different entities in the code snippet and prompt-based approaches to detect entities in summary. We note that detecting entities in the generated summary is more difficult due to the high degree of polysemy Tabassum et al. (2020). For example, entities like "list", "while", "if", etc, can be a code entity or natural language entities. This necessitates our reliance on large language models with high reasoning capabilities for detecting entities on the summary side. We then map the detected entities from the summary to code by using string-matching heuristics. The sentences with unmapped entities can be considered as ungrounded (source of extrinsic hallucination). For each mapped entity, we then have a tuple <code, entity, intent-related sentence>, where the intent-related sentence can be considered as the sentence in summary mentioning the entity. The final step is to verify each tuple from the summary for intrinsic hallucination to assess the correctness of the code summary. Our experiments demonstrate the importance of localizing entities in the summary for effective hallucination detection. Our contributions are :
@@ 2 Related Work
Hallucination in Natural Langauge: Rawte et al. (2024); Sahoo et al. (2024) review recent advances in hallucination detection in natural language, emphasizing its practical significance. Recently, prompt-based methods (Arora et al. (2022), Manakul et al. (2023), Agrawal et al. (2023), Dhuliawala et al. (2023)) are being used to detect hallucinations in the text produced by LLMs. Xiao and Carenini (2022) and Zhang et al. (2022) attempt to address entity-level verification in natural language inputs. Both of these works involve improving the correctness of natural language summaries and do not discuss anything in the context of code. We note that most of the hallucination detection frameworks Manakul et al. (2023); Arora et al. (2022); Dhuliawala et al. (2023); Valentin et al. (2024); Rebedea et al. (2023) in natural language do not enforce reference text for grounding. In our setup of code summarisation, the generated summary has to be evaluated with respect to a reference text (the code snippet). Therefore, neccessiating an approach which could compare the code summary to the code snippet. Maynez et al., 2020; Ji et al., 2023 discuss further fine-graining of hallucination in natural language as intrinsic and extrinsic hallucination. More specifically, Intrinsic hallucination occurs when the given text contradicts the reference, while Extrinsic hallucination happens when the text cannot be verified against the reference. We use a similar convention in our paper.
Hallucination in Code Generation: The code generation space has captured significant attention due to its practical significance in software development. Jiang et al. (2024) discusses recent developments in code generation and suggests the importance of addressing hallucination for improving the reliability of LLMs. Liu et al. (2024) studies hallucination in code generation and proposes a categorization that encompasses five categories of hallucinations based on the conflicting objectives and varying degrees of deviation observed in code generation. Tian et al., 2024; Agarwal et al., 2024; Spracklen et al., 2024 advanced the field with datasets and frameworks addressing hallucination in code generation. These studies highlight that while LLM-generated code may be syntactically correct and semantically plausible, it often fails to execute as intended or meet requirements.
Despite progress in hallucination detection, code summarization remains underexplored. Kang et al., 2024 and Zhang, 2024 focused on inconsistencies in comment generation, addressing specific aspects like design constraints and parameter types, but their methods face challenges due to reliance on execution environments. In contrast, our approach validates the full functionality of generated outputs, independent of external dependencies, offering a more reliable solution by grounding entities in the input code and verifying their intent.
@@ 3 Datasets
To create the hallucination dataset for code summarization, we consider code snippets from Java programming language and CodeXGLUE Lu et al. (2021) – Code-To-Text dataset. We focused on Java programming language due to its widespread relevance in the industry. It offers a rich set of entities (such as classes, methods, and variables) due to its structured design and strict typing system. The dataset was annotated by 8 annotators who are experts in Java and held at least a Master’s degree in Computer Science, with some having a PhD in the field. On average, the annotators had 4+ years of experience in Java programming. We report the statistics in Table 1 and describe the data curation process below:
Summary Generation: We generate summaries from CodeXGLUE by prompting seven different LLMs (Appendix A) with 600 code snippets. By producing multiple summary variants, we can assess hallucination generation by different LLMs and evaluate hallucination detection techniques under varied conditions. We present quantitative results (Table 3) and qualitative analysis in Section 3. During initial annotation, we found that annotators spent considerable time verifying summaries, often requiring online documentation searches, leading to an average annotation time of 30 minutes or more per summary. To ensure feasibility, we randomly prune samples and use $\sim$ 10% of the data for the final hallucination annotation task (Table 1).
@@
Hallucination Labeling: For each detected code entity in a summary, all sentences describing that entity are considered relevant. To account for the scenario where the relevant sentence can be noisy, we introduced a third label, "IRRELEVANT", which can be used to evaluate the performance of the intent-detection module and removed during the preprocessing. Thus, we obtain tuples of (code, entity, relevant sentences) for each entity. A total of 9933 such tuples, sampled from 441 summaries, were selected for human annotation (hired based on volunteering) to detect hallucinations. Out of these, 4354 tuples (from 222 summaries) were independently reviewed by two different sets of annotators, leading to a Cohen Kappa score of 0.72, implying high agreement. The conflicts were resolved by two independent meta-annotators. The annotators were asked to evaluate overall summary by assigning a label of ‘GOOD", ‘FAIR" and ‘POOR" We observe that, on average, 1.33 entities were marked as hallucinated for the summaries rated as ’FAIR" or ’POOR". Therefore, we consider a summary as hallucinated if at least one of the entities is hallucinated. After pre-processing, we consider the instance with labels "CORRECT" and "INCORRECT" in human data and treat the "IRRELEVANT" label predicted by the model as "INCORRECT". We provide the complete annotation guideline in Appendix E.2.
@@ 4 Categorization of Factors for Hallucination in Code Summarization
In this section, we describe the various factors that could lead to hallucination in code summaries (Figure 2) based on what we learned from the annotation process. This classification, based on the underlying factors of hallucination, offers insights into the generative behaviors of language and code models. We describe these categories of hallucination factors below and discuss their statistical analysis in Figure 3.
@@
1) Contextual code:This occurs when the model fails to correctly explain the code, often because it has not encountered the functionality of code during training or is working with a low-resource language like COBOL, where fundamental rules may be misrepresented in the summary.
@@
HC3: Code Complexity: This pertains to the model’s tendency to produce incorrect code summaries due to high code complexity. This may stem from the model’s insufficient reasoning capabilities to understand the code or the misinterpretation of user instructions. Key factors contributing to complexity include: 1) Length: Longer code is more complex, requiring deeper understanding and involving more interdependencies with more potential points of failure. 2) Lexical Complexity: Complex vocabulary, including diverse operands and operations, increases the number of elements to track and understand. 3) Logical Complexity: Code with high cyclomatic complexity, multiple paths, or distant method invocations increases the challenge of comprehension.
@@
HC4: Natural language Context: This refers to cases where natural language in code snippets, such as outdated comments or log statements, causes hallucinations in code summaries. In the code snippet shown in HC4 Example, the LLama3-70B model incorrectly infers that the property variable contains a list of key-value pairs inferred from a commented line. However, the ‘property" variable contains an alphanumeric string followed by one or more semi-colons.
@@ 5.1 Entity Verification
Entity Extraction from summary: Tabassum et al. (2020) propose the task of entity detection in code summaries and introduce a relevant NER tagset. We adopt this tagset for extracting entities from code summaries (Prompt: Appendix A Figure 5).
Leveraging LLMs to recognize entities introduces the risk of hallucinations, where the model may fabricate entities not present in the code summary. To address this, we implement a filtration step to remove such fabricated entities. We evaluate Gemini and GPT-4-Omni for Code NER using human-collected data, with results in Appendix C. Additionally, we assess an open-source model as part of our contribution. Our findings show a strong correlation between GPT-4-Omni predictions and human data, confirming its effectiveness for entity detection in our framework.
@@ 5.2 Entity-Intent Verification
The presence of an entity in both the summary and code indicates that the entity is valid but does not warrant the correctness of the context in which it is discussed. For example, in Figure 1 jobId is a correct entity, but the context of retrieving jobID from the database is incorrect. To address this problem, we propose verifying whether the intent of each mapped entity is accurately described in the summary. We extract all sentences containing the entity of interest from the summary to form its intent context. To identify these relevant sentences that describe an entity’s intent, we explored two approaches: (1) prompting a language model to find relevant sentences and (2) using string-matching heuristics (Refer Appendix D). Qualitative assessment showed that rule-based heuristics were more effective and efficient than prompt-based methods, which suffered from hallucinations. Therefore, we relied on string-matching-heuristics for our framework. After identifying the entity and intent, we use LLMs with zero-shot prompting to verify their correctness with the code (Prompt: Appendix 5).We also experimented with few-shot prompting by including examples of various hallucination types in code summaries along with the representative code. However, performance degraded due to the increased prompt length, consistent with findings in recent works like Mirzadeh et al. (2024).
@@ 5.3 Instance Level Hallucination Detection
Figure 3: Distribution of different hallucination categories proposed in the taxonomy. We observe that the models tend to hallucinate most frequently due to the high complexity of the code, while significant instances of insufficient knowledge were also identified.
@@ 6 Experiments and Results
For our experiments, we consider summaries from instruction-tuned versions of the SOTA code and language models, from IBM-Granite family (20B-instruct; 34B-instruct) Mishra et al. (2024), Llama3 family (8B-instruct and 70B-instruct) Touvron et al. (2023), CodeLlama family (7B and 34B) Roziere et al. (2023) and Mistral family (7B-instruct) Jiang et al. (2023).
For intent verification, we started our initial experiments with powerful open-source models like LLama3-70B Touvron et al. (2023) but observed worse than random performance (F1 score- 0.37) due to the very high complexity of the task. Similar qualitative studies with other models in our preliminary experiments prompted us to explore advanced models with stronger capabilities in code reasoning, such as GPT4-Omni Achiam et al. (2023) and Gemini-1.5-Flash Team et al. (2023). Hyperparameters and other details can be found in Appendix (B).
Model P R F1
@@
Gemini-1.5-Direct 0.41 0.49 0.25
Gemini-1.5-ETF* 0.68 0.64 0.51
GPT4-Omni-Direct 0.48 0.50 0.28
GPT4-Omni-ETF* 0.72 0.74 0.73
@@
Gemini-1.5 0.55 0.70 0.55
GPT4-Omni 0.59 0.69 0.61
Table 2: Hallucination Detection results. We consider two evaluation aspects: 1) Instance Level, which aims to label the entire summary, and 2) Entity level, which labels individual entities in summaries. Here, P, R and F1 refer to the precision, recall and F1 score.
@@ 7.1 Quantitative Analysis
As shown in Table 3, Granite-20B produced shorter summaries, while Llama3-70B generated longer ones. Other models had similar average lengths, reflecting varying elaboration due to differences in training methodologies. For entity mapping, we observe that Llama3-70B has the most mapped entities, indicating the tendency of the model to stay grounded. Granite-20B has the most unmapped entities, which indicates its tendency to produce content which may not be directly related to the code, leading to extrinsic hallucination.
@@ 7.2 Predictive Analysis
Our framework captures ungrounded entities in summaries. For example, when the model refers to a non-existent Hub API based on the keyword ‘HubException", the unmapped entities such as "Hub" are identified. This enables fine-grained and interpretable detection of hallucinations. We can see that the ungrounded entities like "Hub" are captured here, leading to more fine-grained and interpretable detection of hallucination.
[ht!] private List<Transaction> retrieveTransactions(String rowStatusCd) throws HubException {................} Summary: This method retrieves transactions using the “Hub" API based on the input parameters by the “Hub database". It interacts with several key entities, including “squid:s1166" and “squid:s1172", to gather the data…… Analysis: Hallucination and Unmapped Entities